Shift Summary: Owl Shift 6/03/99

Miguel Barrio (mbarrio@hep.uchicago.edu)
Thu, 3 Jun 1999 07:09:43 -0500

Shift Summary for Owl shift, 06/03/99

On Shift: Julie Whitmore
Miguel Barrio

Others present:
Val and Ed working on the CsI MF.
Rick futzed with the filter planes hanging up to no avail until
he discovered a sem file name was incorrect (and cursed quite a bit
about it...)

Hours of beam delivered at all intensities: 7.0
Hours of beam delivered at nominal intensity: 7.0
Hours of 99 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): ?

Accounting of lost time:

Beam was on most of the night, but problems with the DA intereferred
a bit with our runs.

Activities:

0:15 Run 13309 starts, hangs up;
0:40 Run 13310 starts;
0:45 Rick Ford rings us, beam will be delivered soon;
1:45 Run 13311 starts;
2:35 Run 13311 stopped cuz CSQ6 constants are no good;
2:50 Run 13312 starts w/o DC monitoring to see if this is hanging
up da system. Rick discovers that this only affects the
monitoring plane, which is doing fine, the problem is in the
filter planes;
3:15 End of run 13312, install new CSQ6 constants;
3:30 Run 13313 starts;
4:10 Too much shtuffs in home disk, end of 13313 run. Plane1-Stream2
problem is kicking in (CsI data corrupted, only at high rate),
log files are filling up disks. Clean up with `remove_logfiles`.
Problem seems to appear with ECAL config only;
4:15 Run 13314 starts;
4:25 HV for HA's, COSM & CA restored (somehow they got screwed up);
4:35 RC9 trigger source is gone;
5:00 Run 13315 starts;
5:15 Called OPS to get beam stop in for muons;
5:50 Magnet rampdown;
6:15 Run 13321 with muon config starts;
6:30 Run 13322 starts, turning off bananas and TRD monitoring and
turning DC's monitoring back on.
Also, RC9 is alright though the trigger source is dead;
7:00 Ended muon run and ramped the magnet to Ptkick = 411 MeV/c.

=========================================================================
Runs ended this shift:

13309-13310;
13311 cutesie DC time dist's (yeah!), PMT gains r ogly;
13312;
13313 very nice and relatively long run;
13314-13322.

Tapes brought to FCC this shift: 0

-- 
Miguel Barrio
Office: HEP 319 (773-702-7484)
Email:  mbarrio@hep.uchicago.edu